# TMI! Finetuned Models Leak Private Information from their Pretraining Data

### v2

@@ A.1 Introduction
In this mean estimation setting, $\|\nu\|_{2}$ can be thought of as the inversely proportional to the similarity between the pretraining and finetuning tasks. If $\mu$ is similar to the mean of the new data, $Y$, $\|\nu\|_{2}$ is small. The term, $\alpha$, is analagous to the fraction of pretraining epochs (i.e. number of pretraining epochs divided by the total number of pretraining and finetuning epochs). For example if there are 80 pretraining epochs and 20 finetuning epochs, the corresponding $\alpha$ value would be $0.8$. Note that as $\alpha\to 0$, the information from the empirical mean of $X$ is completely overshadowed by the empirical mean of $Y$. Prior work on membership-inference attacks on machine learning models has suggested that gradient updates (a special case of mean estimation) that do not contain an individual make membership-inference success decrease with respect to that individual [37]. This is consistent with the results we present in this section for the simplified setting of membership-inference attacks on Gaussian mean estimation.
@@
[…]
@@ A.3 Results
[…]
@@
[…]
@@ C.1 Datasets
  • CIFAR-100: The CIFAR-100 [39] dataset is a subset of the Tiny Images dataset [5], provided by the Canadian Institute for Advanced Research. It is comprised of 60,000 32x32 color images from 100 classes, where each class contains 600 images (500 for training and 100 for testing). CIFAR-100 is used as one of our pretraining tasks because it is a challenging dataset with a wide variety of classes, which allows our models to learn very general features and patterns that can be applied to several downstream tasks.
@@
  • Tiny ImageNet: Tiny ImageNet [56] is an image classification dataset designed to be a smaller scale alternative to the popular ImageNet [21] dataset. This dataset contains 110,000 64x64 color images from 200 classes, where each class contains 550 images (500 for training and 50 for testing). We use Tiny ImagenNet to pretrain the larger Wide ResNet architecture because it provides a similarly general task to CIFAR-100 at a larger scale. Additionally, the full-sized version of ImageNet is a widely used dataset for pretraining large image models, thus reinforcing the need to evaluate our attack on a dataset like Tiny ImageNet.
@@
  • Coarse CIFAR-100: The classes in CIFAR-100 can be divided into 20 superclasses. Each image in the dataset has a "fine" label to indicate its class and a "coarse" label to indicate its superclass. We construct this coarse dataset using the superclass labels and use it as our downstream task with the highest similarity to the pretraining task. In our experimentation, we ensure that this downstream task does not contain any of the pretraining samples from the standard CIFAR-100 dataset.
@@
  • Caltech 101: Caltech 101 [41] is an image classification dataset comprised of about 9000 color images from 101 classes, where each class contains 40 to 800 images. Because the images in this dataset vary in size and tend to be high resolution, we downscale them to be 64x64 to reduce computational complexity. We use the Caltech 101 dataset to finetune our pretrained Tiny ImageNet models as it provides a difficult task with many categories that can be solved by leveraging the generic features learned during pretraining.
@@
  • Oxford-IIIT Pet: The Oxford-IIIT Pet Dataset [49] is made up of about 7400 color images of cats and dogs. This dataset contains 37 classes with roughly 200 images per class. In our evaluation, this downstream task is the least similar to CIFAR-100 because it focuses on a specific category of images that are mutually exclusive to the pretraining set classes. For this task, we do not use feature extraction because the finetuned model have low utility. Rather, we use the pretrained model as an initialization and update all of its weights.
@@
  • Caltech 101: In a similar fashion to CIFAR-100, the CIFAR-10 [39] is comprised of 60,000 32x32 color images selected from the Tiny Images dataset. This dataset includes 10 classes, each containing 6000 points (5000 for training and 1000 for testing) which are mutually exclusive to those seen in CIFAR-100. In our evaluation, this downstream task is the second most similar to CIFAR-100 because they are both derived from the same distribution of web-scraped images, but they are disjoint in their classes. Although the classes do not overlap, the features learned from pretraining on CIFAR-100 may be useful in performing this task.
@@
  • WikiText-103: WikiText-103 [46] is a large-scale language dataset that is widely used for benchmarking language models. It contains over 100 million tokens and is derived from several Wikipedia articles and contains a vast amount of textual data. The language models we consider in this paper have been pretrained on the train partition of WikiText-103 and are hosted on Hugging Face.
@@
  • DBpedia: The DBpedia ontology (or topic) classification dataset [67] is composed of 630,000 samples with 14 non-overlapping classes from DBpedia, which is a project aiming to extract structured content from the information on Wikipedia. For each of the 14 topics, there are 40,000 training samples and 5000 testing samples. In our experiments with language models, we update a subset of the model’s weights on random subsets of this dataset.
@@
  • Yahoo Answers: The Yahoo Answers topic classification dataset [68] is composed of 1.4 million training samples and 60,000 testing samples with 10 classes. The training and testing sets are divided evenly amongst the topics, such that there are 140,000 training samples and 6000 testing samples per class. Each sample contains both the title and content of a question asked on Yahoo Answers. In our experiments with language models, we update a subset of the model’s weights on random subsets of this dataset, where the content of each question is appended to the question title.
@@ C.2 Shadow Model Training
Our shadow model training procedure for vision models is split into two phases: pretraining and finetuning. In the first phase, we train 129 randomly initialized ResNet models on random subsets of Tiny ImageNet and CIFAR-100, each containing half of the dataset (50k and 25k points, respectively). The remaining samples are held out for evaluation. We train each of the ResNet-34 models for 100 epochs (to 75-80% top-5 accuracy) using SGD with weight decay ($\lambda=10^{-5}$) and cosine annealing [45] as our learning rate scheduler. Using the same hyperparameters, we train the Wide ResNet-101 architecture for 200 epochs to 60% top-5 accuracy. When training and querying any of our shadow models, we use standard data augmentations, such as random crops and horizontal flips.

### v3

@@ A.1 Introduction
In this mean estimation setting, $\|\nu\|_{2}$ can be thought of as the inversely proportional to the similarity between the pretraining and finetuning tasks. If $\mu$ is similar to the mean of the new data, $Y$, $\|\nu\|_{2}$ is small. The term, $\alpha$, is analagous to the fraction of pretraining epochs (i.e. number of pretraining epochs divided by the total number of pretraining and finetuning epochs). For example if there are 80 pretraining epochs and 20 finetuning epochs, the corresponding $\alpha$ value would be $0.8$. Note that as $\alpha\to 0$, the information from the empirical mean of $X$ is completely overshadowed by the empirical mean of $Y$. Prior work on membership-inference attacks on machine learning models has suggested that gradient updates (a special case of mean estimation) that do not contain an individual make membership-inference success decrease with respect to that individual [jagielski-2023-forgetting]. This is consistent with the results we present in this section for the simplified setting of membership-inference attacks on Gaussian mean estimation.
@@
(b) ROC of the membership inference attack on mean estimation for varying amounts of distribution shift, $\|\nu\|_{2}$. In each of these simulations, $\alpha$ is set optimally according to Lemma A.2
@@ A.3 Results
${\mathbb{E}\mathopen{}\mathclose{{}\left(z}\right)}=0\quad\textrm{and}\quad{\mathrm{Var}\mathopen{}\mathclose{{}\left(z}\right)}=d\tilde{\alpha},$
@@
${\mathbb{E}\mathopen{}\mathclose{{}\left(z}\right)}=\frac{\alpha d}{n}\quad\textrm{and}\quad{\mathrm{Var}\mathopen{}\mathclose{{}\left(z}\right)}=d\tilde{\alpha}+\frac{2d\alpha^{2}}{n^{2}}$
@@ C.1 Datasets
  • CIFAR-100: The CIFAR-100 [CIFAR] dataset is a subset of the Tiny Images dataset [TinyImages], provided by the Canadian Institute for Advanced Research. It is comprised of 60,000 32x32 color images from 100 classes, where each class contains 600 images (500 for training and 100 for testing). CIFAR-100 is used as one of our pretraining tasks because it is a challenging dataset with a wide variety of classes, which allows our models to learn very general features and patterns that can be applied to several downstream tasks.
@@
  • Tiny ImageNet: Tiny ImageNet [tiny-imagenet] is an image classification dataset designed to be a smaller scale alternative to the popular ImageNet [imagenet] dataset. This dataset contains 110,000 64x64 color images from 200 classes, where each class contains 550 images (500 for training and 50 for testing). We use Tiny ImagenNet to pretrain the larger Wide ResNet architecture because it provides a similarly general task to CIFAR-100 at a larger scale. Additionally, the full-sized version of ImageNet is a widely used dataset for pretraining large image models, thus reinforcing the need to evaluate our attack on a dataset like Tiny ImageNet.
@@
  • Coarse CIFAR-100: The classes in CIFAR-100 can be divided into 20 superclasses. Each image in the dataset has a ”fine” label to indicate its class and a ”coarse” label to indicate its superclass. We construct this coarse dataset using the superclass labels and use it as our downstream task with the highest similarity to the pretraining task. In our experimentation, we ensure that this downstream task does not contain any of the pretraining samples from the standard CIFAR-100 dataset.
@@
  • Caltech 101: Caltech 101 [caltech-101] is an image classification dataset comprised of about 9000 color images from 101 classes, where each class contains 40 to 800 images. Because the images in this dataset vary in size and tend to be high resolution, we downscale them to be 64x64 to reduce computational complexity. We use the Caltech 101 dataset to finetune our pretrained Tiny ImageNet models as it provides a difficult task with many categories that can be solved by leveraging the generic features learned during pretraining.
@@
  • Oxford-IIIT Pet: The Oxford-IIIT Pet Dataset [pets] is made up of about 7400 color images of cats and dogs. This dataset contains 37 classes with roughly 200 images per class. In our evaluation, this downstream task is the least similar to CIFAR-100 because it focuses on a specific category of images that are mutually exclusive to the pretraining set classes. For this task, we do not use feature extraction because the finetuned model have low utility. Rather, we use the pretrained model as an initialization and update all of its weights.
@@
  • Caltech 101: In a similar fashion to CIFAR-100, the CIFAR-10 [CIFAR] is comprised of 60,000 32x32 color images selected from the Tiny Images dataset. This dataset includes 10 classes, each containing 6000 points (5000 for training and 1000 for testing) which are mutually exclusive to those seen in CIFAR-100. In our evaluation, this downstream task is the second most similar to CIFAR-100 because they are both derived from the same distribution of web-scraped images, but they are disjoint in their classes. Although the classes do not overlap, the features learned from pretraining on CIFAR-100 may be useful in performing this task.
@@
  • WikiText-103: WikiText-103 [wikitext103] is a large-scale language dataset that is widely used for benchmarking language models. It contains over 100 million tokens and is derived from several Wikipedia articles and contains a vast amount of textual data. The language models we consider in this paper have been pretrained on the train partition of WikiText-103 and are hosted on Hugging Face.
@@
  • DBpedia: The DBpedia ontology (or topic) classification dataset [dbpedia] is composed of 630,000 samples with 14 non-overlapping classes from DBpedia, which is a project aiming to extract structured content from the information on Wikipedia. For each of the 14 topics, there are 40,000 training samples and 5000 testing samples. In our experiments with language models, we update a subset of the model’s weights on random subsets of this dataset.
@@
  • Yahoo Answers: The Yahoo Answers topic classification dataset [2015-yahoo] is composed of 1.4 million training samples and 60,000 testing samples with 10 classes. The training and testing sets are divided evenly amongst the topics, such that there are 140,000 training samples and 6000 testing samples per class. Each sample contains both the title and content of a question asked on Yahoo Answers. In our experiments with language models, we update a subset of the model’s weights on random subsets of this dataset, where the content of each question is appended to the question title.
@@ C.2 Shadow Model Training
Our shadow model training procedure for vision models is split into two phases: pretraining and finetuning. In the first phase, we train 129 randomly initialized ResNet models on random subsets of Tiny ImageNet and CIFAR-100, each containing half of the dataset (50k and 25k points, respectively). The remaining samples are held out for evaluation. We train each of the ResNet-34 models for 100 epochs (to 75-80% top-5 accuracy) using SGD with weight decay ($\lambda=10^{-5}$) and cosine annealing [loshilov-2017-annealing] as our learning rate scheduler. Using the same hyperparameters, we train the Wide ResNet-101 architecture for 200 epochs to 60% top-5 accuracy. When training and querying any of our shadow models, we use standard data augmentations, such as random crops and horizontal flips.